OpenStudy (anonymous):

how to do this kind of question 6x2+19x+10=0

hero (hero):

Hint: \(15 + 4 = 19 \\ 15 \times 4 = 60\)

OpenStudy (anonymous):

still cant get it

hero (hero):

@greenlegodude57, it's \(6x^2 + 19x + 10 = 0\)

OpenStudy (anonymous):

Ohhhh, sorry I read the question wrong.

OpenStudy (anonymous):

@ashiq What are the choices?

hero (hero):

Replace 19 with 15 + 4: \(6x^2 + (15 + 4)x + 10 = 0\) Then: \(6x^2 + 15x + 4x + 10 = 0\) Now factor by grouping

OpenStudy (anonymous):

thx got it

hero (hero):

What do you mean you got it? What did you get?

OpenStudy (anonymous):

by the way hows 4 come

OpenStudy (anonymous):

the soloution to solve the question